A carrom board (40cm)(40cm) has queen at the centre.When the queen is hit by the striker, it moves to the front edge, rebounds and goes into the hole behind the striking line. The displacement if the queen is

Grade:6

1 Answers

Prsahant Parashar
7 years ago
Displacement is only from mid point of hypotaneous to the corner 
 
 
length of hypotaneous=(40+402)1/2 = 40.21/2
 
displacement =56.56/2 =28.28